# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ 7 Conclusion
We address the challenges of both data heterogeneity and model heterogeneity in PFL. While prior works typically tackle only one of these challenges, or both under simplified assumptions, we propose FedMosaic, which handles both forms of heterogeneity comprising two key components: SIMA and PQ-LoRA. Moreover, to better reflect real-world data heterogeneity in PFL scenarios, we introduce DRAKE, a comprehensive multi-modal FL benchmark that captures both task heterogeneity and distribution shifts. Our extensive empirical evaluation across diverse PFL scenarios shows that FedMosaic not only achieves superior personalization, outperforming local training, but also improves generalization and few-shot adaptation, which are essential for future personalization.
Acknowledgment
We acknowledge support from the European Research Council (ERC-2020-AdG 101021347 KeepOnLearning). This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT): No.RS-2022-II220077, No.RS-2022-II220113, No.RS-2022-II220959, No.RS-2022-II220871, No.RS-2022-II220951 (50%), No.RS-2021-II211343 (SNU AI), No.RS-2021-II212068 (AI Innovation Hub), and IITP-2025-RS-2024-00437633 (ITRC). Additional support was provided by the BK21 FOUR program, SNU in 2025.
@@ Appendix A Technical Appendices and Supplementary Material
  • Section A.10: Additional experiments using Qwen-based LLaVA models.
  • Section A.11: Results on the Fed-LLM benchmark, complementing our Fed-Aya experiments.
@@
  • Section A.13: Analysis of how the low rank parameter $r$ in PQ-LoRA affects performance.
  • Section A.14: Ablation study on SIMA, comparing different strategies for measuring task similarity.
@@
  • Section A.16: Study on the effect of varying the number of layers employing PQ-LoRA.
Benchmark Details and Limitations
  • Section A.18: Comprehensive description of our proposed DRAKE benchmark, including task categories, dataset sources, and client configuration.
@@ A.1 Theoretical Justification for Freezing $A$ and $B$ in PQ-LoRA
We first define the aggregation error $\delta$. In an FL setup with $N$ clients, the server aggregates client updates $\Delta W_{i}=B_{i}P_{i}A_{i}$ for $i\in[N]$. With FedAvg aggregation, the ‘ideal’ aggregation $\Delta W^{*}$ (sun2024improving, ; guo2025selective, ) should be:
@@
However, we cannot perform ‘ideal’ aggregation in FL, since $W^{*}$ cannot be decomposed into trainable parameters, i.e., $B$, $P$, and $A$, on client sides (guo2025selective, ). In other words, although $W^{*}$ can be computed on the server, it cannot be redistributed as trainable parameters on the client side. As a result, instead of directly averaging $\Delta W_{i}$ for $i\in[N]$, we average the trainable parameters separately and obtain the practical update $\Delta\tilde{W}^{*}$ as:
@@
To this end, we define the aggregation error $\delta=|\Delta W^{*}-\Delta\tilde{W}^{*}|$. During local training, we update shareable modules $P$ and $Q$, while freezing $A$ and $B$ to preserve alignment. This way, we minimize the aggregation error $\delta$, following Theorem 2.
@@
If $A$ is frozen after alignment, then $A=A_{1}=A_{2}=\cdots=A_{N}$. Similarly, freezing $B$ implies $B=B_{1}=B_{2}=\cdots=B_{N}$. Under this condition, both the ideal update $\Delta W^{*}$ (Eq. 11) and the practical update $\Delta\tilde{W}^{*}$ (Eq. 12) can be simplified to:
@@ A.2 Proof of Theorem 1
To satisfy Eq.19, it is sufficient for the column vectors of $B$ and the row vectors of $A$ to form linearly independent sets, even if they are not strictly orthogonal. However, compared to initialization with linearly independent vectors, orthogonal initialization offers additional advantages, such as faster (hu2020provable, ) and stable (nowak24a, ) convergence. Therefore, motivated by both the practical advantages and the theoretical justification provided in Theorem 1, we initialize $A$ and $B$ with orthogonal sets.
@@ A.3 Details of Client Model Selection in the Heterogeneous PFL Scenario
In the heterogeneous PFL scenarios, we assign each client’s model based on supervised fine-tuning (SFT) performance on their respective tasks. Specifically, clients are assigned larger models when these models exhibit significantly better SFT performance for their tasks. Conversely, when the performance difference between model sizes is negligible, we allocate smaller models to maximize computational efficiency, as larger models require substantially more computational resources. We summarize the SFT performance of each LLaVA-Llama3 variant model and the resulting client-wise model allocation in Tab. 6.
Model DRAKE
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
Model HFLB
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
Table 6: Per-client SFT performance of different LLaVA-Llama3 model variants
A.4 More Details of PQ-LoRA
@@
However, during the alignment of $A_{i}$ and $A_{j}$ in Eq. 8, orthogonality can be disrupted due to MSE loss training. To prevent significant deviation from orthogonality, we add a regularization term to the objective function of $A$ alignment as follows:
$\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\mathrm{MSE}\bigl{(}A_{i}(x),A_{j}(x)\bigr{)}+\lambda\left\|A_{j}^{\top}A_{j}-I\right\|_{F}^{2},$ (22)
@@
A.5 Details of Experimental Setups
@@
To employ LLaVA-Llama3 variant models (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, LLaVA-Llama3.1-8B) and LLaVA-Qwen2.5 variant models (i.e., LLaVA-Qwen2.5-0.5B, LLaVA-Qwen2.5-1.5B, and LLaVA-Qwen2.5-3B) in heterogeneous PFL scenarios, we instruction-tune them on LLaVA-Instruct-158K (liu2023visual, ), following the original LLaVA training setup (liu2023visual, ). We summarize the performance of the instruction-tuned variants in Tab. 7.
Model VQAv2 GQA VizWiz SciQA TextVQA POPE MME MMBench SEED LLaVA MM-Vet
(goyal2017vqav2, ) (hudson2019gqa, ) (gurari2018vizwiz, ) (lu2022sciqa, ) (singh2019textvqa, ) (li2023pope, ) (fu2024mme, ) (liu2024mmbench, ) -Bench (li2023seedbench, ) -Wild (liu2023visual, ) (yu2023mmvet, )
LLaVA-1.5-7B (liu2023visual, ) 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.6 65.4 31.1
LLaVA-Llama3.2-1B 75.3 59.7 45.3 67.3 49.0 85.0 1338.1 61.3 57.8 59.6 28.7
LLaVA-Llama3.2-3B 78.7 62.6 45.3 75.71 55.8 86.1 1446.5 70.9 62.5 67.0 34.5
LLaVA-Llama3.1-8B 79.8 64.2 47.6 76.4 57.9 86.3 1486.1 72.5 64.1 73.7 32.7
LLaVA-Qwen2.5-0.5B 71.8 56.5 37.3 58.1 43.4 86.2 1251.9 53.6 53.1 55.0 23.1
LLaVA-Qwen2.5-1.5B 75.3 59.3 40.9 69.5 51.8 86.4 1376.4 66.8 60.6 60.0 27.9
LLaVA-Qwen2.5-3B 77.2 61.4 50.9 74.1 54.7 87.2 1447.6 71.4 63.2 66.7 33.1
Table 7: Zero-Shot performance of instruction-tuned LLaVA-1.5 variants. We use the CLIP ViT-L/336px model as the vision encoder for all variants.
@@
We evaluate FedMosaic on Multi-modal FL benchmarks, such as our proposed DRAKE and HFLB, and LLM-FL benchmarks, such as Fed-LLM and Fed-Aya. We summarize the details of benchmarks in Tab. 8. As shown in the table, unlike HFLB, which focuses solely on VQA tasks, DRAKE covers diverse task types and involves more images per sample due to its multi-image inputs. In DRAKE-Dynamic, each client learns four tasks, where all data samples are randomly mixed in PFL-Static, while the samples from each task are introduced incrementally in PFL-Dynamic. We provide more detail of DRAKE in Sec. A.18.
HFLB includes 7 VQA datasets, where we assign them into 9 clients by partitioning GQA hudson2019gqa into three clients based on the QA type (Yes/No, 2 Multi-choice, and 4 Multi-choice). We further split each assigned dataset into four tasks to allow experiments on the PFL-Dynamic setup. We use keywords in questions or GPT to categorize each sample, where the details are provided in Tab. 9.
Benchmark ##\## of Images ##\## of Questions ##\## of clients ##\## of rounds $R$#\## of local steps Task types
@@
HFLB (chen2024feddat, ) 217K 314K 9 20 100 VQA
Fed-LLM (ye2024fedllm, ) - 30K 5 20 30 General Capability, Mathematics, Coding
Fed-Aya (singh2024aya, ) - 36K 8 20 50 Question-Answering in 12 Languages
@@
Client 1 Client 2 Client 3 Client 4 Client 5
GQA (hudson2019gqa, ): Yes/No
@@
GQA (hudson2019gqa, ): Four-choice
@@
Abstract VQA (antol2015abstract, )
@@
Client 6 Client 7 Client 8 Client 9
COCOQA (ren2015cocoqa, )
@@
NLVR2 (suhr-etal-2019nlvr, )
@@
We measured the accuracy metrics, $A_{last}$ and $A_{\text{AUC}}$, based on the correct choice for multi-choice questions and the correct tokens compared to the ground-truth answer for the open-ended questions. For Fed-Aya, we used GPT to rate the generated response compared to the given ground-truth response. We use the same prompt template shown in Fig. 7 from the original paper (singh2024aya, ). For Fed-LLM, we follow the evaluation process of the original paper (ye2024fedllm, ) and use MMLU (hendrycks2020mmlu, ), GSM8K (cobbe2021gsm8k, ), and HumanEval (chen2021humaneval, ) benchmarks to evaluate general capability, math and coding skills, respectively.
@@
We set the batch size to 4, the learning rate for PQ-LoRA to $5\times 10^{-5}$, and for other parameters to $2\times 10^{-5}$. We use the Constant LR scheduler and the AdamW optimizer (loshchilov2019decoupled, ) for all datasets. We use the LoRA rank $r$ of 128 attached to all linear layers in the LLM backbone. For LLM experiments, we set the learning rate to $3\times 10^{-4}$ and the LoRA rank to 16. For federated learning, we set the total number of communication rounds $R$ to 20. For a fair comparison, we adjust the local training step for each baseline and FedMosaic to use the same computational cost as SFT, following  (seo2025budgeted, ). For the PFL-Dynamic setup, we employ memory-only training, where newly encountered samples are added to episodic memory, and training batches are retrieved solely from memory, following  (koh2023online, ; seo2024learning, ; seo2025budgeted, ). For episodic memory, we adopt a memory-infinite setup (prabhu2023online, ; seo2025budgeted, ), assuming all data can be stored in memory, reflecting that memory cost is not a bottleneck in real-world scenarios. For $\tau$, the softmax temperature parameter used in SIMA, we use 0.5 in all experiments. For $\lambda$, loss balancing coefficient between MSE loss and regularization term in Eq. 22, we use 0.5. We set $N_{B}$, the number of layers employing PQ-LoRA, to 4 in all experiments.
@@
For $W_{s}$, a small-scale pre-trained MLLM used for calculating client-wise gradients in SIMA, we adopt the smallest model among all client models. For example, in Table 11, where clients have three different types of heterogeneous architectures (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B), we employ LLaVA-Llama3.2-1B as $W_{s}$ for all clients. This choice ensures computational efficiency and maintains gradient dimension consistency for cosine similarity calculation. All experiments are executed in Python 3.10, on four Ubuntu 20.04 machines, with 8 NVIDIA RTX A6000 GPUs each. Each experiment runs on a single RTX A6000 GPU in a day.
A.6 Comparison of Memory and Computational Costs
We compare the computational cost $\mathcal{C}$ and memory cost $\mathcal{M}$ of various baselines and summarize the results in Tab. 10. Following (seo2025budgeted, ), we measure computational cost in FLOPs per iteration and memory cost in Bytes. Specifically, for each baseline, we report the relative FLOPs and relative Bytes in comparison to supervised fine-tuning (SFT), which only requires a single forward and backward pass without any extra computation and memory overhead.
Comparison of Computational Cost $\mathcal{C}$.
We first compare the computational cost of FL methods. PerAda (xie2024perada, ) incurs approximately twice the computational cost compared to other baselines, since it sequentially updates both the personalized adapter and the local adapter. Similarly, FedDAT (chen2024feddat, ) independently optimizes the local adapter and the dual adapter teacher (DAT), which also results in a double computational cost. Our proposed FedMosaic incurs a minor additional computational cost due to gradient computation from the frozen pre-trained model used for measuring task similarity and dual adapter structure, but this overhead amounts to only approximately 9.8% more FLOPs. The reason for the small increase in computation is that we only use the last layer gradient of the small-scale pre-trained frozen model for computational efficiency, as mentioned in Sec. 4.1. Specifically, we minimize gradient computation overhead by using only the last-layer gradients of the frozen LLaVA-1.5-1B model on one batch out of every 10 iterations, rather than computing gradients on all data. Moreover, note that while FedMosaic also introduces a dual adapter structure, i.e., maintaining the local adapter and the global adapter separately, similar to PerAda, FedDAT, and FedDPA, we optimize the dual-adapter at once using a learnable balancing parameter in PQ-LoRA, which adaptively balances local and global output. In contrast, the baselines separately optimize each adapter, thus incurring a double computational cost.
Comparison of Memory Cost $\mathcal{M}$.
We then compare the memory cost of FL methods. DITTO, FedSim, PerAda, FedDAT, FedDPA, as well as FedMosaic employ a dual adapter structure, which maintains both local and global adapters separately, to preserve global knowledge. It incurs additional memory cost, but since we adopt the LoRA adapter (hu2022lora, ), which occupies significantly less memory compared to pre-trained weights (qi2024fdlora, ), the dual adapter only consumes about 5% additional memory cost compared to using a single LoRA adapter.
Methods Memory Cost $\mathcal{M}$ Computational Cost $\mathcal{C}$
Overhead Type Relative $\mathcal{M}$ to SFT Overhead Type Relative $\mathcal{C}$ to SFT

### v3

@@ 7 Conclusion
We address both data heterogeneity and model heterogeneity in PFL while prior works tackle only one challenge or both under simplified assumptions. For that, we propose FedMosaic, which handles both forms of heterogeneity through two components: RELA and PQ-LoRA. Moreover, to better reflect real-world data heterogeneity in PFL scenarios, we introduce DRAKE, a comprehensive multi-modal FL benchmark capturing task heterogeneity and distribution shifts. Extensive evaluation shows FedMosaic achieves superior personalization compared to local training, while improving generalizability and few-shot adaptation capabilities essential for future personalization.
Ethics Statement
We propose a better learning scheme for heterogeneous federated learning for realistic federated learning scenarios. While the authors do not explicitly aim for this, the increasing adoption of deep learning models in real-world contexts with streaming data could potentially raise concerns, such as inadvertently introducing biases or discrimination. We note that we are committed to implementing all feasible precautions to avert such consequences, as they are unequivocally contrary to our intentions.
Reproducibility statement
We take reproducibility in deep learning very seriously and highlight some of the contents of the manuscript that might help to reproduce our work. We will definitely release our implementation of the proposed method in Sec. 4, benchmark in Sec. 5, and the data splits and the baselines used in our experiments in Sec. 6.
The use of Large Language Models
We use large language models (LLMs) to support labor-intensive and mistake-prone work. Specifically, we use LLMs (e.g., GPT-4) to categorize samples from a single dataset into multiple tasks based on keywords or topics, to assess generation quality in specific benchmarks using LLMs, and to detect grammatical errors during writing.
@@ Appendix A Technical Appendices and Supplementary Material
  • Section A.4: Consistency of FedMosaic over diverse random heterogeneous PFL scenarios.
@@
  • Section A.7: Summary of client model configuration of all heterogeneous PFL experiments.
@@
  • Section A.9: Comparative analysis of communication (transmission) cost of FedMosaic.
@@
  • Section A.13: Experimental results ins heterogeneous PFL scenarios using Qwen-based LLaVA models.
  • Section A.14: Additional experiments on various cross-family heterogeneous PFL setups.
  • Section A.15: Results on the text-only benchmark, complementing our Fed-LLM-Large experiments.
@@
  • Section A.17: Analysis of the choice (i.e., type and size) of public data $D_{P}$ for PQ-LoRA alignment.
  • Section A.18: Analysis of how the change of various hyperparameters introduced in FedMosaic affects performance.
  • Section A.19: Ablation study on RELA, comparing different strategies for measuring task similarity.
  • Section A.20: Analysis on the size of $W_{s}$ for RELA in terms of computational cost and performance in a PFL setup.
@@
  • Section A.22: Discussion of RELA and existing similarity-based model aggregation in federated learning and multi-task learning.
Limitations/Future Work and Benchmark Details
  • Section A.24: Comprehensive description of our proposed DRAKE benchmark, including task categories, dataset sources, client configuration, and comparison with existing FL benchmarks.
  • Section A.25: Comparison between DRAKE and existing FL benchmarks.
@@ A.1 Theoretical Justification for Freezing $A$ and $B$ in PQ-LoRA
We first define the aggregation error $\delta$. In an FL setup with $N$ clients, the server aggregates client updates $\Delta W_{i}=B_{i}P_{i}A_{i}$ for $i\in[N]$. With FedAvg aggregation, the ‘ideal’ aggregation $\Delta W^{*}$ (Sun et al., 2024; Guo et al., 2025) should be:
@@
However, we cannot perform ‘ideal’ aggregation in FL, since $W^{*}$ cannot be decomposed into trainable parameters, i.e., $B$, $P$, and $A$, on client sides (Guo et al., 2025). In other words, although $W^{*}$ can be computed on the server, it cannot be redistributed as trainable parameters on the client side. As a result, instead of directly averaging $\Delta W_{i}$ for $i\in[N]$, we average the trainable parameters separately and obtain the practical update $\Delta\tilde{W}^{*}$ as:
@@
[…]
@@
If AA is frozen after alignment, then $A=A_{1}=A_{2}=\cdots=A_{N}$. Similarly, freezing $B$ implies $B=B_{1}=B_{2}=\cdots=B_{N}$. Under this condition, both the ideal update $\Delta W^{*}$ (Eq. 9) and the practical update $\Delta\tilde{W}^{*}$ (Eq. 10) can be simplified to:
@@ A.2 Proof of Theorem 1
To satisfy Eq.17, it is sufficient for the column vectors of $B$ and the row vectors of $A$ to form linearly independent sets, even if they are not strictly orthogonal. However, compared to initialization with linearly independent vectors, orthogonal initialization offers additional advantages, such as faster (Hu et al., 2020) and stable (Nowak et al., 2024) convergence. Therefore, motivated by both the practical advantages and the theoretical justification provided in Theorem 1, we initialize $A$ and $B$ with orthogonal sets.
@@ A.3 Details of Client Model Selection in the Heterogeneous PFL Scenario
In the heterogeneous PFL scenarios, we assign each client’s model based on supervised fine-tuning (SFT) performance on their respective tasks. Specifically, clients are assigned larger models when these models exhibit significantly better SFT performance than smaller ones. Conversely, when the performance difference between model sizes is negligible, we allocate smaller models to maximize computational efficiency, as larger models require substantially more computational resources. We summarize the SFT performance of each LLaVA-Llama3 variant model and the resulting client-wise model allocation in Tab. 8.
DRAKE
Model Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
HFLB
Model Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
Table 8: Per-client SFT performance of different LLaVA-Llama3 model variants. We assign the larger model (i.e., LLaVA-Llama-3.2-3B) to clients whose SFT performance shows a substantial gain, while the smaller model (i.e., LLaVA-Llama-3.2-1B) is assigned otherwise.
A.4 Effect of Client Model Assignment
To validate whether FedMosaic consistently achieves strong performance regardless of the specific model assigned to each client, we conduct experiments with varying ratios of LLaVA-1B and LLaVA-3B models, where model assignments are randomly determined in each run. In Table 9, we summarize the PFL results of supervised fine-tuning (SFT) and FedMosaic across diverse heterogeneous PFL scenarios, each corresponding to a different configuration of client model selections in the DRAKE-dynamic benchmark. We observe that FedMosaic consistently outperforms SFT with minimal performance variance across heterogeneous setups, demonstrating its robustness to the composition of client models.
Self Others
Scenario Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
$\Delta$ (FedMosaic- SFT) +1.45 +1.39 +3.75 +2.82
[…]
$\Delta$ (FedMosaic- SFT) +2.20 +1.86 +4.38 +3.02
[…]
$\Delta$ (FedMosaic- SFT) +1.62 +1.64 +4.38 +3.08
[…]
$\Delta$ (FedMosaic- SFT) +2.26 +1.94 +3.94 +3.02
[…]
$\Delta$ (FedMosaic- SFT) +2.68 +2.02 +3.87 +2.71
[…]
A.5 More Details of PQ-LoRA
@@
However, during the alignment of $A_{i}$ and $A_{j}$ in Sec. 4.2.2, orthogonality can be disrupted due to L2 loss training. To prevent significant deviation from orthogonality, we add a regularization term to the objective function of $A$ alignment as follows:
$\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\|(A_{i}(x)-A_{j}(x)\|^{2}_{2}+\lambda\left\|A_{j}^{\top}A_{j}-I\right\|_{F}^{2},$ (20)
@@
A.6 Details of Experimental Setups
@@
To employ LLaVA-Llama3 variant models (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, LLaVA-Llama3.1-8B) and LLaVA-Qwen2.5 variant models (i.e., LLaVA-Qwen2.5-0.5B, LLaVA-Qwen2.5-1.5B, and LLaVA-Qwen2.5-3B) in heterogeneous PFL scenarios, we instruction-tune them on LLaVA-Instruct-158K (Liu et al., 2023b), following the original LLaVA training setup (Liu et al., 2023b). We summarize the performance of the instruction-tuned variants in Tab. 10 and Tab. 11. We use Llama-3 base models for text-only benchmarks.
VQAv2 GQA VizWiz SciQA TextVQA
Model (Goyal et al., 2017) (Hudson & Manning, 2019) (Gurari et al., 2018) (Lu et al., 2022) (Singh et al., 2019)
LLaVA-1.5-7B 78.5 62.0 50.0 66.8 58.2
LLaVA-Llama3.2-1B 75.3 59.7 45.3 67.3 49.0
LLaVA-Llama3.2-3B 78.7 62.6 45.3 75.7 55.8
LLaVA-Llama3.1-8B 79.8 64.2 47.6 76.4 57.9
LLaVA-Qwen2.5-0.5B 71.8 56.5 37.3 58.1 43.4
LLaVA-Qwen2.5-1.5B 75.3 59.3 40.9 69.5 51.8
LLaVA-Qwen2.5-3B 77.2 61.4 50.9 74.1 54.7
Table 10: Zero-Shot performance of instruction-tuned LLaVA-1.5 variants on academic-task-oriented benchmarks. We use the CLIP ViT-L/336px model as the vision encoder for all variants.
POPE MME MMBench SEED-Bench LLaVA-Wild $M$-Vet
Model (Li et al., 2023c) (Fu et al., 2023a) (Liu et al., 2024) (Li et al., 2023a) (Liu et al., 2023b) (Yu et al., 2023)
LLaVA-1.5-7B 85.9 1510.7 64.3 58.6 65.4 31.1
LLaVA-Llama3.2-1B 85.0 1338.1 61.3 57.8 59.6 28.7
LLaVA-Llama3.2-3B 86.1 1446.5 70.9 62.5 67.0 34.5
LLaVA-Llama3.1-8B 86.3 1486.1 72.5 64.1 73.7 32.7
LLaVA-Qwen2.5-0.5B 86.2 1251.9 53.6 53.1 55.0 23.1
LLaVA-Qwen2.5-1.5B 86.4 1376.4 66.8 60.6 60.0 27.9
LLaVA-Qwen2.5-3B 87.2 1447.6 71.4 63.2 66.7 33.1
Table 11: Zero-Shot performance of instruction-tuned LLaVA-1.5 variants on instruction-following benchmarks. We use the CLIP ViT-L/336px model as the vision encoder for all variants.
@@
We evaluate FedMosaic on Multi-modal FL benchmarks, such as our proposed DRAKE and HFLB, and LLM-FL benchmarks, such as Fed-Scope and Fed-Aya, and Fed-LLM-Large. We summarize the details of benchmarks in Tab. 12. As shown in the table, unlike HFLB, which focuses solely on VQA tasks, DRAKE covers diverse task types and involves more images per sample due to its multi-image inputs. In DRAKE-Dynamic, each client learns four tasks, where all data samples are randomly mixed in PFL-Static, while the samples from each task are introduced incrementally in PFL-Dynamic. We provide more details of DRAKE in Sec. A.24.
HFLB includes 7 VQA datasets, where we assign them into 9 clients by partitioning GQA Hudson & Manning (2019) into three clients based on the QA types (e.g., Yes/No, 2 Multi-choice, and 4 Multi-choice). We further split each assigned dataset into four tasks to allow experiments on the PFL-Dynamic setup. We use keywords in questions or GPT to categorize each sample, where the details are provided in Tab. 13.
To simulate PFL in NLP, we modify existing sets (i.e., Fed-Aya and Fed-Scope) and compose larger mixtures (i.e., Fed-LLM-Large) to increase task variety and heterogeneity across clients. Specifically, Fed-Aya is composed of 12 different languages equally sampled from different language families. We further separate the tasks based on the topic of the QA using GPT-4. Fed-Scope (Kuang et al., 2024) includes coding, mathematics, and general capability datasets, where each client learns either one of the three datasets. Fed-LLM-Large is a large-scale text-only benchmark designed for personalized federated learning. To diversify the personal tasks among clients, we combined tasks from three sources, Fed-ChatbotIT and Fed-aya from Fed-LLM (Ye et al., 2024), and Fed-FLAN Long et al. (2024). This results in 52 clients and 2 tasks per client, where each client fine-tunes on either instruction following tasks from different sources or in different languages (note that all other benchmarks have 4 tasks per client).
Benchmark #\# of Images #\# of Questions #\# of clients #\# of rounds $R$ #\# of local steps Task types
@@
HFLB 217K 314K 9 20 100 VQA
Fed-Scope - 30K 5 20 30 General Capability, Mathematics, Coding
Fed-Aya - 36K 8 20 50 Question-Answering in 12 Languages
Fed-LLM-Large - 31K 52 4 10 Instruction Following
@@
Client 1 Client 2 Client 3
GQA (Hudson & Manning, 2019)
: Yes/No
@@
GQA (Hudson & Manning, 2019)
: Four-choice
@@
Client 4 Client 5 Client 6
Abstract VQA (Antol et al., 2015)
@@
COCOQA (Ren et al., 2015)
@@
Client 7 Client 8 Client 9
NLVR2 (Suhr et al., 2019)
@@
We measured the accuracy metrics, $A_{last}$ and $A_{\text{AUC}}$, based on the correct choice for multi-choice questions and the correct tokens compared to the ground-truth answer for the open-ended questions. For Fed-Aya, we used GPT to rate the generated response compared to the given ground-truth response. We use the same prompt template shown in Fig. 8 from the original paper (Singh et al., 2024). For Fed-Scope, we follow the evaluation process of the original paper (Kuang et al., 2024) and use MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), and HumanEval (Chen et al., 2021) benchmarks to evaluate general capability, math and coding skills, respectively. For Fed-LLM-Large, we use Rouge-L (Lin, 2004) metric as it is commonly used metric to assess LLM generation quality compared to ground-truth (Long et al., 2024; Li et al., 2024c).
@@
We set the batch size to 4, the learning rate for PQ-LoRA to $5\times 10^{-5}$, and for other parameters to $2\times 10^{-5}$. We use the Constant LR scheduler and the AdamW optimizer (Loshchilov & Hutter, 2019) for all datasets. We use the LoRA rank $r$ of 128 attached to all linear layers in the LLM backbone. For LLM experiments, we set the learning rate to $3\times 10^{-4}$ and the LoRA rank to 16. For federated learning, we set the total number of communication rounds $R$ to 20, with the local training step of 100. For a fair comparison, we adjust the local training step for each baseline and FedMosaic to use the same computational cost as SFT, following  (Seo et al., 2025). For the PFL-Dynamic setup, we employ memory-only training, where newly encountered samples are added to episodic memory, and training batches are retrieved solely from memory, following  (Koh et al., 2023; Seo et al., 2024; 2025). For episodic memory, we adopt a memory-infinite setup (Prabhu et al., 2023; Seo et al., 2025), assuming all data can be stored in memory, reflecting that memory cost is not a bottleneck in real-world scenarios. For $\tau$, the softmax temperature parameter used in RELA, we use 0.5 in all experiments. For $\alpha$, the EMA ratio of client-wise gradient in Eq. 2, we use 0.5 in all experiments. For $\lambda$, the loss balancing coefficient between the L2 loss and the regularization term in Eq. 20, we use 0.5. For $N_{s}$, gradient sampling ratio for sanitized gradient $\tilde{g_{i}}$, we use 40%. For $\mu$, the noise scale applied to $\epsilon$, we set $\mu=10^{-4}$. We set $N_{B}$, the number of layers employing PQ-LoRA, to 4 in all experiments. See Sec. A.18 for detailed hyperparameter analysis.
@@
Federated distillation baselines (i.e., TAKFL, PerAda, and FedMKT) share knowledge between heterogeneous models via logits, while FedMosaic employs PQ-LoRA. For the other baselines, we combine them with Fed-ET (Cho et al., 2022), which aggregates models among homogeneous clients. For $W_{s}$, a small-scale pre-trained MLLM used for calculating client-wise gradients in RELA, we adopt the smallest model among all client models. For example, in Table 16, where clients have three different types of heterogeneous architectures (i.e., LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B), we employ LLaVA-Llama3.2-1B as $W_{s}$ for all clients. This choice ensures computational efficiency and maintains gradient dimension consistency for cosine similarity calculation. Note that we can employ a lighter $W_{s}$ (e.g., reduced size or lower-bit quantization) to further decrease the computation overhead of gradient calculation in RELA, as detailed in Sec. A.20. All experiments are executed in Python 3.10, on four Ubuntu 20.04 machines, with 8 NVIDIA RTX A6000 GPUs each. Each experiment runs on a single RTX A6000 GPU in a day.
A.7 Client Model Configurations
We summarize the model configurations for each PFL experiment, including model types and their counts, in Tab. 14. For experiments on multi-modal benchmarks, we use LLaVA-Llama3 or LLaVA-Qwen2.5 variants, while for text-only benchmarks, we use Llama-3 variants.
LLaVA-Llama3 LLaVA-Qwen2.5 Llama-3
Benchmark Experiment 1B 3B 8B 0.5B 1.5B 3B 1B 3B 8B
Multi-modal Tab. 2 DRAKE-Dynamic 4 6 0 0 0 0
Tab. 2 HFLB-Dynamic 3 6 0 0 0 0
Tab. 4 DRAKE-Static 4 6 0 0 0 0
Tab. 4 DRAKE-Dynamic 0 3 0 0 4 3
Tab. 7 DRAKE-Dynamic 4 6 0 0 0 0
Tab. 16 DRAKE 3 5 2 0 0 0
Tab. 17 DRAKE-Homo 0 10 0 0 0 0
Tab. 17 HFLB-Homo 0 9 0 0 0 0
Tab. 18 DRAKE-Homo 0 0 0 0 0 10
Tab. 18 DRAKE-Hetero 0 0 0 3 2 5
Tab. 19 Llama 3B / Qwen 1.5B 0 6 0 0 4 0
Tab. 19 Llama 1B / 3B / Qwen 1.5B 2 5 0 0 3 0
All other analysis / ablation experiments 4 6 0 0 0 0
Text-only Tab. 6 Fed-LLM-Large-Dynamic 26 26 0
Tab. 7 Fed-Scope-Static 0 3 2
Tab. 20 Fed-aya-Dynamic 4 4 0
Tab. 20 Fed-Scope-Static 0 3 2
Table 14: Client model configuration details for each experiment. Counts per model family/size.
A.8 Comparison of Computational and Memory Costs
Memory Cost $\mathcal{M}$ Computational Cost $\mathcal{C}$
Methods Overhead Type Relative $\mathcal{M}$ to SFT Overhead Type Relative $\mathcal{C}$ to SFT